# Inflationary Krylov complexity

### 17 Jan 2024 v1

@@ Abstract
In this work, we have systematically investigated the Krylov complexity of the modified dispersion relation in inflation, using the algorithm in closed system and open system. Since many quantum gravitational frameworks could lead to this kind of modified dispersion relation, our analysis could be applied to the string cosmology, loop gravity, $\it e.t.c$. Following the Lanczos algorithm, we find the very early universe is an infinite, many-body, and maximal chaotic system. Our numerics shows that the Lanczos coefficient and Lyapunov index of the standard dispersion relation are mainly determined by the scale factor. As for the modified case, it is nearly determined by the momentum. In a method of the closed system, we discover that the Krylov complexity will show irregular oscillation before the horizon exits. The modified case will present faster growth after the horizon exists. As for the approach of an open system, we construct the exact wave function which is very robust only requiring the Lanczos coefficient proportional to $n$ (main quantum number). Based on it, we find the Krylov complexity and Krylov entropy could nicely recover in the case of a closed system under the weak dissipative approximation, in which our analysis shows that the evolution of Krylov complexity will not be the same with the original situation. Meanwhile, our numerics clearly shows the Krylov complexity will grow during the whole inflationary period. But for the small scales, there will be a peak after the horizon exits. Our analysis reveals that the dramatic change in background (inflation) will significantly impact the evolution of Krylov complexity. Since the curvature perturbation will transit from the classical level to the quantum level. We could expect that the decoherence will highly impact the Krylov complexity during inflation.
@@
The holographic principle needs the conformal symmetry for the quantum field theory. Meanwhile, with the development of complexity in high-energy physics, it is natural to investigate the complexity in quantum systems. Ref. Aaronson:2016vto defined complexity as the minimal operations for a certain task, namely also dubbed as the computational complexity. Afterward, there were two main methods for calculating the complexity: (a) The geometric method of Nielsen et al Nielsen2006AGA ; Nielsen2006QuantumCA ; Dowling2008TheGO . (b) The method of ”Fubini-Study” distance on the related quantum states Chapman:2017rqy . To be more precise, these two main methods, especially the Nielson’s geometric method, are aimed to investigate the so-called circuit complexity, what they investigate is the minimal number of quantum operations from a reference state to a target state. The third method was called the Krylov complexity, Ref. Parker:2018yvk introduced the Krylov complexity as the measurement of operator growth with respect to Krylov basis, in which they proposed a universal operator growth in the chaotic system. Compared with the previous two methods, the definition of Krylov complexity is free of ambiguities since its definition is unique. For the other two methods, the complexity depends on the choice of quantum gates (the geometric method) or the metric (the method of ”Fubini-Study” distance). Once the quantum state is defined, one can construct the Krylov space in light of the Lanczos algorithm viswanath1994recursio , and one can calculate the Krylov complexity, Lanczos coefficient, and its corresponding entropy. Ref. Muck:2022xfc systematically developed the method for calculating the Krylov complexity in various orthogonal polynomials dubbed as the inner product. The Krylov complexity can be applied for the SKY model Parker:2018yvk ; Rabinovici:2020ryf ; Jian:2020qpp . One can also define a thermalized Krylov complexity Kar:2021nbm . Even the chaos can be understood as the delocalization in Krylov space Dymarsky:2019elm . By contrast, Ref. Rabinovici:2021qqt proposed the Krylov complexity is suppressed by the finite size which is in the integrable models. For further tests and numerical calculations, it can be implemented into the generalized coherent states Caputa:2021sib ; Patramanis:2021lkx , the Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , CFT Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye , integrable models with saddle-point dominated scrambling Bhattacharjee:2022vlt , cosmological complexity Adhikari:2022oxr $\it e.t.c$. The very recent developments can be found in Refs. Erdmenger:2023wjg ; Patramanis:2023cwz ; Fan:2023ohh ; Hashimoto:2023swv ; Vasli:2023syq ; Domingo:2023kjr ; Gill:2023umm ; Bhattacharjee:2023uwx .
@@
One of the main motivations is that we aim to explore the Krylov complexity in the very early universe since the energy of the universe is not conserved due to the symmetry broken along the temporal part. It is naturally considering our universe as an open system. To be more precise, one could consider that the universe will exchange information with the environment (outer of the universe). Following Refs. Bhattacharya:2023zqt ; Bhattacharjee:2022lzy , we will write the Hamilton of single field inflationary model with modified dispersion relation into one part that governs the closed and unitary system, another part will response the information exchanging with the environment. This kind of treatment can be considered as the Markovian environment in terms of Lindbladian evolution Lindblad:1975ef ; Gorini:1975nb . It should be noticed that the whole Hamilton is non-Hermitian under this Lindbladian evolution. From Ref. Socorro:2022aoz , it indicates that the gravity system (k-essence) is non-Hermitian. Therefore, it is quite natural to use the approach of open system.
@@
We organize the paper as follows. In Sec. II, we will give a brief introduction to Krylov complexity and the Lanczos algorithm, which is the method for dealing with closed system. In Sec. III.1, some fundamental knowledge about the modified dispersion relation and inflation will be given to readers. Sec. IV will give the calculation of Krylov complexity and its corresponding Krylov entropy and discuss some properties in inflation. In Sec. V, we will generalize the Lanczos algorithm into the open system and discuss the general properties of Krylov complexity in a weak dissipative system. Then, we will construct an exact wave function for calculating the Krylov complexity and Krylov entropy. Finally, the conclusion and outlook will be given in Sec. VI.
@@ III.2 Basics of inflation
IV Krylov complexity of modified dispersion relation in closed system
@@ IV.2 Krylov complexity, Lanczos coefficient and Lyapunov index
V Krylov complexity in open system
In this section, we will implement the method of Bhattacharya:2022gbz to investigate the Krylov complexity in an open system, which the very early universe is dubbed as an open system.
@@ V.2 General discussions of $\phi(n)$
where we have made the approximations: $b_{n+1}\approx b_{n}$ and $\phi_{n+1}-\phi_{n-1}\approx 2\partial_{n}\phi$. Following Ref. Bhattacharjee:2022lzy , its contious version can be written by $\partial_{\eta}\phi+n[\chi\mu\phi+2\alpha\partial_{n}\phi]=0$. Making a comparison, one can obtain $\tilde{c}_{n}=n\chi\mu$ and $b_{n}=n\alpha$. As for $\eta$ approaches to infinitite, Eq. (44) will be of the form $\phi(n)\propto e^{-n/\xi}$ with $\xi=\frac{2\alpha}{\chi\mu}$. Since $\mu$ presents the strength between the dynamical system and the environment, one can consider that the dynamical system is nearly a closed system as $\mu$ approaches zero and vice versa. Accordingly, $\xi$ will be very large as $\mu$ is very tiny. The corresponding evolution of complexity in the weak dissipation region will be in this way: At the very beginning of dynamical system, the Krylov complexity will spread exponentially whose average position is $K(\eta)\propto e^{2\alpha\eta}$ as $\eta<\eta_{*}$, which it is still true until $K(\eta)$ saturates to $\xi$ and $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln(\frac{2\alpha}{\chi\mu})$. It is straightforward to see that the $\eta_{*}$ is just the $\xi$ nothing else. As for the strong dissipation region, the growth of Krylov complexity is different.
@@
V.3 Wave function of $\phi(n)$ in open system
@@
where $C$ is a coefficient. Compared with general discussion of $\phi$, we could recover the factor $e^{-\frac{n}{\xi}}$ and we could find the missing factor $n^{\frac{in\omega}{2b_{n}}}$. Based on the general discussion of V.2, we can use the results to analyze our situation: (a) The Krylov complexity will be exponentially growing with $e^{2\alpha\eta}$ with
$\alpha=\sqrt{(k/2)^{2}(f^{2}-1)^{2}+(a^{\prime}/a)^{2}},$ (54)
where the conformal time $\eta$ is less than $\eta_{*}$. (b) As $\eta\textgreater\eta_{*}$, the Krylov complexity will saturate to
@@
Noticing that $\xi$ in our case is not a constant anymore, its varying trend depends on $k$ and $a$, which will lead to that $\eta_{*}$ is also varying since $\eta_{*}$ is just $\xi$ as discussed in Sec. V.2. For a better understanding of the evolution of Krylov complexity, we plot it with $\eta\textgreater\eta_{*}$ and $\eta\textless\eta_{*}$.
Figure 4: The plot of Krylov complexity with $\eta\textgreater\eta_{*}$ and $\eta\textless\eta_{*}$. $y=\log_{10}(a)$ varies from -4 to 0. The ”horizon exit” is located at $y=-2$ (still in the very early universe). We have set $k=0.01$ (ensuring the weak dissipative region) and $H_{0}=1$.
Fig. 4 clearly shows the evolution of Krylov complexity with two cases: $\eta\textgreater\eta_{*}$ and $\eta\textless\eta_{*}$, where the range of $y$ will confirm the weak dissipative region: (a) As $\eta<\eta_{*}$, the Krylov complexity of standard case will almost keep a constant value. But for the modified cases, it shows the Krylov complexity will sharply increase to a peak and then decrease to almost zero, finally, it will exponentially grow until it is constant. This peak appears after the horizon exits. Thus, one could see that the appearance of a peak could be dubbed as a criterion for assessing whether it is a modified case or not. (b) As $\eta>\eta_{*}$, one could see that the trend is almost constant until it will exponentially grow as $y>-1$.
One could see that the total trend of Krylov complexity is almost consistent with the general discussion of V.2 in the weak region. However, there is the peak of Krylov entropy as $\eta<\eta_{*}$ which is a new finding in our case. Through observation, one could see that this peak appears after the horizon exits. It is reasonably assumed that this kind of decoherene will cause this kind of peak. Another reason may come from the geometrical effects of the background.
We will make some comments on the evolution of Krylov complexity. First and foremost, the general discussion is valid for the system whose background is static. Once taking the inflationary effects into account, we will find that the evolution of complexity will show various evolutions, even if the varying trend is totally different at small scales. Another fact is that the various quantum gravitational frameworks will also lead to the varying trend of Krylov complexity. Especially, the Krylov complexity will decrease for the modified dispersion relation as $\alpha\textgreater 1$ which was not discovered before. Finally, we only plot the evolution of some factors of the Krylov complexity since it is not from the complete wave function, which is shown in Sec. V.2. To completely understand the evolution of Krylov complexity, we will need the exact wave function of $\phi(n)$.
@@ V.4 Exact wave function of $\phi(n)$ in an open inflationary universe
This wave function is our central result in this paper which describes the opening system including our inflationary universe. Based on it, we will implement to calculate the Krylov complexity and Krylov entropy in weak and strong dissipative region.
First, we will investigate the weak dissipative region. As $\eta$ approaches infinity, the momentum will decrease dramatically due to the exponential expansion of the universe, thus one can see that $u_{1}$ and $u_{2}$ will be much smaller than unity. Another way is that the wave function (V.4) will behave like
@@
where we have set $n=1$. Doing the Taylor expansion of $u_{1}^{2}$ and $u_{2}$ of wave function (59), one can obtain $\phi_{n}(\eta\rightarrow\infty)\propto 1-u_{2}-\frac{1}{2}u_{1}^{2}$. According to the general discussion of V.2, the wave function of $\phi_{n}$ should have the form as $\exp(-\frac{n}{\xi})$, its Taylor expansion for the firs order is $1-\frac{1}{\xi}=1-u_{2}-\frac{1}{2}u_{1}^{2}u_{2}$. As $u_{1}\ll 1$ and $u_{2}\ll 1$, thus we could see that the asymptotic wave function could match the stationary solution $\phi_{n}\propto e^{-n/\xi}$. Thus, one can also derive that the $\xi^{-1}\approx u_{2}$ in the weak dispassion region so as to obtain $\eta_{*}=\ln(1/u_{2})/2$.
@@
it clearly shows that the Krylov complexity will be recovered into the form of $\sinh^{2}\eta$, where the Krylov complexity is $\sinh^{2}(r_{k})$ in a closed system whose wave function (32) comes via the two-mode squeezed state. The higher order will contribute to the correction due to the dissipation effects. However, all of these calculations hold true under the assumption with $u_{1},~{}u_{2}\ll 1$. The practical situation is different with this analysis since the $u_{1}$ and $u_{2}$ will be larger than unity, where their values are determined by $u_{1}^{2}=|1-\frac{k^{2}}{4}(f^{2}-1)^{2}-\frac{a^{\prime 2}}{a^{2}}|$ and $u_{2}^{2}=\frac{k^{2}}{4}(f^{2}+1)^{2}$. We will show the exact evolution of Krylov complexity according to (61).
@@ V.6 Krylov entropy in an open inflationary universe
where we have used $-\rm sech^{2}\eta/A^{2}=-\cosh^{2}\eta+4u_{2}\cosh^{3}\eta\sinh\eta$ expanding around zero of $u_{2}$ and we only retain the first order of $u_{1}$ or $u_{2}$. For the leading order, it is consistent with the formula of (39) which confirms the wave function (59) is correct. The correction that is proportional to $u_{2}$ comes via the dispersion effects.
@@
Krylov complexity could provide a unique definition for calculating the complexity, which is free of ambiguities compared with the traditional computational complexity. With the application of quantum information theory, the Krylov complexity will be more and more important in high-energy physics. From another aspect, inflation naturally gives a platform for high energy physics, such as the generation of particles, $\it e.t.c$, where the most important place is that the background of the universe will experience an exponential expansion within a very short temporal interval that may lead to the difference of Krylov complexity. The original method for calculating the Krylov complexity is based on static and flat spacetime. How to develop the method for calculating the Krylov complexity for open system is of particular interest. In this paper, we focus on the framework of modified dispersion relation in inflation. Since many quantum gravitational frameworks could lead to this kind of modified dispersion relation, our analysis is naturally applicable to them, such as the string cosmology, loop gravity, Lorentz violation, and so on.
@@
$(c).$ In the approach of an open system, the investigation is based on our central result, an exact wave function (59) describing the open inflationary universe. Compared with the original method for constructing the wave function Bhattacharjee:2022lzy , our wave function (59) could describe the strong and weak dissipative region. In the weak dissipative approximation, the calculation for Krylov complex (62) and Krylov entropy (64) could nicely recover into the case of a closed system, which ensures the correctness of the wave function. In the weak dissipative region, our calculation could return into the analysis of Bhattacharjee:2022lzy , where there is scrambling time dubbed as a boundary for studying the complexity. The scrambling time is just $\xi$ (55) which is a variable determined by $k$ and $a$. Consequently, it will lead to the Krylov complexity different in the weak dissipative region as shown in Fig. 3. As $\eta<\eta_{*}$, there is a peak of complexity at some small scales which is our new findings as $\eta<\eta_{*}$. In the original discussion of Bhattacharjee:2022lzy , the Krylov complexity will grow exponentially. As for $\eta>\eta_{*}$, the Krylov complexity will grow up to some constant values via Bhattacharjee:2022lzy . Our case is different, it will be always a constant and then it will grow exponentially. Our analysis shows that inflation could show this varying and distinctive trend. Since this peak appears after the horizon exits, it is reasonably assumed that this process may change the trend of Krylov complexity. For the strong dissipative region, one could see that the varying trend is always growing from the large scales, which will not reach some constant values. Fig. 5 clearly indicates the evolution trend of Krylov complexity. Generally, one cannot distinguish the standard case and modified case from the strong dissipative region. At the small scale, there is also one peak of Krylov complexity. One can see that inflation will have a huge impact on the Krylov complexity.
@@
Firstly, our numerics has shown that varying trend of Krylov complexity is enhancing at the whole inflation, which is not consistent with Bhattacharjee:2022lzy , where they show that Krylov complexity will saturate to some constant values. How to confront their results, one possible solution is extending our analysis into the MD and MD. Since the scale factor will evolve differently in these two periods. The various scale factors may cause the Krylov complexity to be saturated to some constant values. Another possibility comes via the decoherence of the curvature perturbations. Ref. Bhattacharjee:2022lzy has investigated the decoherence of the circuit complexity will saturate the constant values, which may be valid for our case.
Secondly, all of our analysis is based on the Hamilton (23) that is Hermitian. The original method for evaluating the Krylov complexity in an open system is based on the Lindblad master equation Lindblad:1975ef ; Gorini:1975nb . The corresponding Lindblad of various dynamical systems is non-Hermitian. For obtaining the non-Hermitian Hamilton, one could expand the action into the higher order in terms of the Mukhanov variable. Another way is that we could add the potential of inflaton in MD and RD. The contribution of potential cannot be ignored due to the violation of slow-roll condition.

### 20 Jan 2024 v2

@@ Abstract
In this work, we have systematically investigated the Krylov complexity of curvature perturbation for the modified dispersion relation in inflation, using the algorithm in closed system and open system. Since many quantum gravitational frameworks could lead to this kind of modified dispersion relation, our analysis could be applied to the string cosmology, loop gravity, $\it e.t.c$. Following the Lanczos algorithm, we find the very early universe is an infinite, many-body, and maximal chaotic system. Our numerics shows that the Lanczos coefficient and Lyapunov index of the standard dispersion relation are mainly determined by the scale factor. As for the modified case, it is nearly determined by the momentum. In a method of the closed system, we discover that the Krylov complexity will show irregular oscillation before the horizon exits. The modified case will present faster growth after the horizon exists. As for the approach of an open system, we construct the exact wave function which is very robust only requiring the Lanczos coefficient proportional to $n$ (main quantum number). Based on it, we find the Krylov complexity and Krylov entropy could nicely recover in the case of a closed system under the weak dissipative approximation, in which our analysis shows that the evolution of Krylov complexity will not be the same with the original situation. We also find the inflationary period is a strong dissipative system. Meanwhile, our numerics clearly shows the Krylov complexity will grow during the whole inflationary period. But for the small scales, there will be a peak after the horizon exits. Our analysis reveals that the dramatic change in background (inflation) will significantly impact the evolution of Krylov complexity. Since the curvature perturbation will transit from the quantum level to the classical level. We could expect that the decoherence will highly impact the Krylov complexity during inflation.
@@
The holographic principle needs the conformal symmetry for the quantum field theory. Meanwhile, with the development of complexity in high-energy physics, it is natural to investigate the complexity in quantum systems. Ref. Aaronson:2016vto defined complexity as the minimal operations for a certain task, namely also dubbed as the computational complexity. Afterward, there were two main methods for calculating the complexity: (a) The geometric method of Nielsen et al Nielsen2006AGA ; Nielsen2006QuantumCA ; Dowling2008TheGO . (b) The method of ”Fubini-Study” distance on the related quantum states Chapman:2017rqy . To be more precise, these two main methods, especially the Nielson’s geometric method, are aimed to investigate the so-called circuit complexity, what they investigate is the minimal number of quantum operations from a reference state to a target state. The third method was called the Krylov complexity, Ref. Parker:2018yvk introduced the Krylov complexity as the measurement of operator growth with respect to Krylov basis, in which they proposed a universal operator growth in the chaotic system. Compared with the previous two methods, the definition of Krylov complexity is free of ambiguities since its definition is unique. For the other two methods, the complexity depends on the choice of quantum gates (the geometric method) or the metric (the method of ”Fubini-Study” distance). Once the quantum state is defined, one can construct the Krylov space in light of the Lanczos algorithm viswanath1994recursio , and one can calculate the Krylov complexity, Lanczos coefficient, and its corresponding entropy. Ref. Aguilar-Gutierrez:2023nyk has shown that one cannot make the Krylov and Nielsen notions of complexity compatible with each other, whereas under certain assumptions, one can allow for Krylov and ”Fuibini-Study” length to be proportional to each other Caputa:2021sib .
Ref. Muck:2022xfc systematically developed the method for calculating the Krylov complexity in various orthogonal polynomials dubbed as the inner product. The Krylov complexity can be applied for the SKY model Parker:2018yvk ; Rabinovici:2020ryf ; Jian:2020qpp . One can also define a thermalized Krylov complexity Kar:2021nbm . Even the chaos can be understood as the delocalization in Krylov space Dymarsky:2019elm . By contrast, Ref. Rabinovici:2021qqt proposed the Krylov complexity is suppressed by the finite size which is in the integrable models. For further tests and numerical calculations, it can be implemented into the generalized coherent states Caputa:2021sib ; Patramanis:2021lkx , the Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , CFT Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye , integrable models with saddle-point dominated scrambling Bhattacharjee:2022vlt , cosmological complexity Adhikari:2022oxr $\it e.t.c$. The very recent developments can be found in Refs. Erdmenger:2023wjg ; Patramanis:2023cwz ; Fan:2023ohh ; Hashimoto:2023swv ; Vasli:2023syq ; Domingo:2023kjr ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf . Especially, Ref. Camargo:2023eev involves examining the Krylov operator complexity within the simple quantum billiard model, which can be considered as the extension to finite temperature of Hashimoto:2023swv . Ref. Huh:2023jxt also has verified all the universal features of spread complexity within saddle dominated scrambling (”integrable” systems).
@@
One of the main motivations is that we aim to explore the Krylov complexity in the very early universe since the energy of the universe is not conserved due to the symmetry broken along the temporal part. It is naturally considering our universe as an open system. There are some attempts for using the complexity in cosmology, see Refs Adhikari:2021ked ; Choudhury:2020hil . To be more precise, one could consider that the universe will exchange information with the environment (outer of the universe). Following Refs. Bhattacharya:2023zqt ; Bhattacharjee:2022lzy , we will write the Hamilton of single field inflationary model with modified dispersion relation into one part that governs the closed and unitary system, another part will response the information exchanging with the environment. This kind of treatment can be considered as the Markovian environment in terms of Lindbladian evolution Lindblad:1975ef ; Gorini:1975nb . It should be noticed that the whole Hamilton is non-Hermitian under this Lindbladian evolution. From Ref. Socorro:2022aoz , it indicates that the gravity system (k-essence) is non-Hermitian. Therefore, it is quite natural to use the approach of open system.
@@
We organize the paper as follows. In Sec. II, we will give a brief introduction to Krylov complexity and the Lanczos algorithm, which is the method for dealing with closed system. In Sec. III, some fundamental knowledge about the modified dispersion relation and inflation will be given to readers. Sec. IV will give the calculation of Krylov complexity and its corresponding Krylov entropy and discuss some properties in inflation. In Sec. V, we will generalize the Lanczos algorithm into the open system and discuss the general properties of Krylov complexity in a weak dissipative system. Then, we will construct an exact wave function for calculating the Krylov complexity and Krylov entropy. Finally, the conclusion and outlook will be given in Sec. VI.
@@ III.2 Basics of inflation
IV Krylov complexity of modified dispersion relation in closed system’s approach
@@ IV.2 Krylov complexity, Lanczos coefficient and Lyapunov index
V Krylov complexity with approach of open system
In this section, we will implement the method of Bhattacharya:2022gbz to investigate the Krylov complexity in an open system’s method, which the very early universe is dubbed as an open system.
@@ V.2 General discussions of $\phi(n)$
where we have made the approximations: $b_{n+1}\approx b_{n}$ and $\phi_{n+1}-\phi_{n-1}\approx 2\partial_{n}\phi$. Following Ref. Bhattacharjee:2022lzy , its contious version can be written by $\partial_{\eta}\phi+n[\chi\mu\phi+2\alpha\partial_{n}\phi]=0$. Making a comparison, one can obtain $\tilde{c}_{n}=n\chi\mu$ and $b_{n}=n\alpha$. As for $\eta$ approaches to infinity, Eq. (44) will be of the form $\phi(n)\propto e^{-n/\xi}$ with $\xi=\frac{2\alpha}{\chi\mu}$. Since $\mu$ presents the strength between the dynamical system and the environment, one can consider that the dynamical system is nearly a closed system as $\mu$ approaches zero and vice versa. Accordingly, $\xi$ will be very large as $\mu$ is very tiny. The corresponding evolution of complexity in the weak dissipation region will be in this way: At the very beginning of dynamical system, the Krylov complexity will spread exponentially whose average position is $K(\eta)\propto e^{2\alpha\eta}$ as $\eta<\eta_{*}$, which it is still true until $K(\eta)$ saturates to $\xi$ and $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln(\frac{2\alpha}{\chi\mu})$. It is straightforward to see that the $\eta_{*}$ is just the $\xi$ nothing else. As for the strong dissipation region, the growth of Krylov complexity is different.
@@
V.3 Wave function of $\phi(n)$ in a weak dissipative system
@@
where $C$ is a coefficient with
@@
and we have obtained the known factor $e^{-\frac{n}{\xi}}$. Until present, we have obtained the solution without any assumption or approximation. Recalling that $\eta$ is just a parameter that does not represent the conformal time since we obtain it only acoording to (53). A careful reader could notice that $\eta$ cannot be larger than zero in inflation in light of $\eta=\frac{-1}{aH}$ where the range of $\eta$ is $-\infty<\eta<0$. It seems that is conflicted with $\eta$ approaches infinity. The truth is that we only consider inflation where the formula of the scale factor is determined. If it is extending our analysis into the late universe, the value of $\eta$ can be the infinity, such as $\eta\propto t^{\alpha}$ ($t$ is the physical time and $\alpha>0$). Consequently, one can conclude that the inflationary period is a strong dissipative system. If one expects to analyze the weak dissipative region, we must extend into the late universe. Thus, the construction of an exact wave function for describing inflation is quite necessary. The function of weak dissipative approximation is that we could test the validity of our exact wave function. Since all of the analysis should be recovered into the case of a closed system under this approximation. In the next section, we will construct an exact wave function for describing a strong dissipative system.
@@ V.4 Exact wave function of $\phi(n)$ in an open inflationary universe
This wave function is our central result in this paper which could desribe the strong and weak dissipative system, of course including the inflation.
First, we need to check that the exact wave function can restore into the case of a closed system under the weak dissipative system, but it is after inflation. As $\eta$ approaches infinity, the momentum will decrease dramatically due to the exponential expansion of the universe, thus one can see that $u_{1}$ and $u_{2}$ will be much smaller than unity. Another way is that the wave function (58) will behave like
@@
where we have set $n=1$. Doing the Taylor expansion of $u_{1}^{2}$ and $u_{2}$ of wave function (58), one can obtain $\phi_{n}(\eta\rightarrow\infty)\propto 1-u_{2}-\frac{1}{2}u_{1}^{2}$. According to the general discussion of V.2, the wave function of $\phi_{n}$ should have the form as $\exp(-\frac{n}{\xi})$, its Taylor expansion for the firs order is $1-\frac{1}{\xi}=1-u_{2}-\frac{1}{2}u_{1}^{2}u_{2}$. As $u_{1}\ll 1$ and $u_{2}\ll 1$, thus we could see that the asymptotic wave function could match the stationary solution $\phi_{n}\propto e^{-n/\xi}$. Thus, one can also derive that the $\xi^{-1}\approx u_{2}$ in the weak dispassion region so as to obtain $\eta_{*}=\ln(1/u_{2})/2$. Until present, our discussion is valid for the weak dissipative system.
@@
it clearly shows that the Krylov complexity will be recovered into the form of $\sinh^{2}\eta$, where the Krylov complexity is $\sinh^{2}(r_{k})$ in a closed system whose wave function (32) comes via the two-mode squeezed state. The higher order will contribute to the correction due to the dissipation effects. However, all of these calculations hold true under the assumption with $u_{1},~{}u_{2}\ll 1$ which furtherly tests the validity of the exact wave function (58). Following the previous discussion, the inflation is a strong dissipative system, we will show the numerics of Krylov complexity (61).
@@ V.6 Krylov entropy in an open inflationary universe
where we have used $-\rm sech^{2}\eta/A^{2}=-\cosh^{2}\eta+4u_{2}\cosh^{3}\eta\sinh\eta$ expanding around zero of $u_{2}$ and we only retain the first order of $u_{1}$ or $u_{2}$. For the leading order, it is consistent with the formula of (39) which confirms the wave function (58) is correct. The correction that is proportional to $u_{2}$ comes via the dissipation effects.
@@
Krylov complexity could provide a unique definition for calculating the complexity, which is free of ambiguities compared with the traditional computational complexity. With the application of quantum information theory, the Krylov complexity will be more and more important in high-energy physics. From another aspect, inflation naturally gives a platform for high energy physics, such as the generation of particles, $\it e.t.c$, where the most important place is that the background of the universe will experience an exponential expansion within a very short temporal interval that may lead to the difference of Krylov complexity. The original method for calculating the Krylov complexity is based on static and flat spacetime. How to develop the method for calculating the Krylov complexity in inflation is of particular interest. In this paper, we focus on the framework of modified dispersion relation in inflation. Since many quantum gravitational frameworks could lead to this kind of modified dispersion relation, our analysis is naturally applicable to them, such as the string cosmology, loop gravity, Lorentz violation, and so on.
@@
$(c).$ In the approach of an open system, the investigation is based on our central result, an exact wave function (58) describing the strong and weak dissipative system, including inflation. Compared with the original method for constructing the wave function Bhattacharjee:2022lzy , our wave function (58) could describe the strong and weak dissipative region. According to $\eta=-\frac{1}{aH}$, one can see that $\eta$ cannot be larger than zero in inflation. But the situation will be different after inflation since the $\eta\propto t^{\alpha}$ in the late universe. Thus, one can conclude that the inflation is a strong dissipative system. However, the analysis can be dubbed as a criterion for testing the validity of the exact wave function (58). As a further test, the calculation for Krylov complex (61) and Krylov entropy (63) could nicely recover into the case of a closed system. In the weak dissipative region, our calculation could return into the analysis of Bhattacharjee:2022lzy , where there is scrambling time dubbed as a boundary for studying the complexity. The scrambling time is just $\xi$ (54) which is a variable determined by $k$ and $a$. For the strong dissipative region, one could see that the varying trend is always growing from the large scales, which will not reach some constant values. Fig. 4 clearly indicates the evolution trend of Krylov complexity. Generally, one cannot distinguish the standard case and modified case from the strong dissipative region. At the small scale, there is also one peak of Krylov complexity. One can see that inflation will have a huge impact on the Krylov complexity.
@@
Firstly, our numerics has shown that varying trend of Krylov complexity is enhancing at the whole inflation, which is not consistent with Bhattacharjee:2022lzy , where they show that Krylov complexity will saturate to some constant values. How to confront their results, one possible solution is extending our analysis into the MD or RD. Since the scale factor will evolve differently in these two periods. The various scale factors may cause the Krylov complexity to be saturated to some constant values. Another possibility comes via the decoherence of the curvature perturbations. Ref. Bhattacharjee:2022lzy has investigated the decoherence of the circuit complexity will saturate the constant values, which may be valid for our case.
Secondly, all of our analysis is based on the Hamilton (23) that is Hermitian. The original method for evaluating the Krylov complexity in an open system is based on the Lindblad master equation Lindblad:1975ef ; Gorini:1975nb . The corresponding Lindblad of various dynamical systems is non-Hermitian. For obtaining the non-Hermitian Hamilton, one could expand the action into the higher order in terms of the Mukhanov variable. Another way is that we could add the potential of inflaton in MD or RD. The contribution of potential cannot be ignored due to the violation of slow-roll condition.